What is the new equation of the conic
x^2 +y^2 +4x−2y+3 = 0, when
i) the origin is shifted at (2,−1), followed by a rotation of axes through 45^◦
?
ii) the axes are rotated through 45^â—¦
, followed by the shifting of the origin at
(2,−1)?
1
Expert's answer
2020-05-11T17:21:39-0400

Given, the equation of conic is "x^2+y^2+4x-2y+3=0 \: (\clubs)" .

i). Now, suppose the conics axes are rotated by "45^{\circ}" and let us consider new axes are "x' \&y'" ,thus

"\begin{bmatrix}
 x' \\
 y'
\end{bmatrix}=\begin{bmatrix}
 \cos(45^{\circ}) & \sin(45^{\circ}) \\
 -\sin(45^{\circ})&\cos(45^{\circ})
\end{bmatrix}\begin{bmatrix}
 x\\
 y
\end{bmatrix} \implies \begin{bmatrix}
 x \\
 y
\end{bmatrix}=\begin{bmatrix}
 \cos(45^{\circ}) & -\sin(45^{\circ}) \\
 \sin(45^{\circ})&\cos(45^{\circ})
\end{bmatrix}\begin{bmatrix}
 x'\\
 y'
\end{bmatrix}" ,

moreover

"x=\frac{x'-y'}{\sqrt{2}} \: (\star)\\ y=\frac{x'+y'}{\sqrt{2}} \: (\star \star)"

On solving the equation "(\clubs)" ,we get,

new equation of conic is "(\frac{x'-y'}{\sqrt{2}})^2+(\frac{x'+y'}{\sqrt{2}})^2 +4(\frac{x'-y'}{\sqrt{2}})-2(\frac{x'+y'}{\sqrt{2}})+3=0" , On simplification equation of conic becomes,

"\sqrt{2}(x'^2+y'^2)+(2x' -6y')+3\sqrt{2}=0 \: (\dag)"

Now , if origin is shifted by

"(2,-1) \implies x'=x''-2 \: \& \: y'=y''+1 \: (\diamond)"

,hence from the "(\dag) \: \& (\diamond)" we get "\sqrt{2}\{(x'' -2)^2+(y''+1)^2)\}+2(x''-2) -6(y'+1')+3\sqrt{2}=0" ,thus on simplification we get the final equation of conic is

"\sqrt{2}(x^2+y^2)+(2-4\sqrt{2})x+(2\sqrt{2} -6)y+8\sqrt{2}=0"


As, "x'',x' \dots y' \&x,y" are dummy variable.



ii). Now, if we shift the origin first and then rotate the axis then exactly similar to above part from"(\diamond) \: \& \: (\clubs)"

if we replace the dummy variable "x'' \leftrightarrow x' \: \& y'' \leftrightarrow y'" ,we get, "\big(\square )"


"(x'-2)^2+(y'+1)^2 +4(x'-2)-2(y'+1) +3=0\\
\implies x'^2+y'^2-2=0"

Now rotate the axes as part (i) equation "(\star), (\star \star)" and considering "\big(\square)" , thus

"(\frac{x''-y''}{\sqrt{2}})^2+(\frac{x''+y''}{\sqrt{2}})^2 -2=0\\
\implies x''^2+y''^2-2=0"

Therefore, the final equation of the conic is

"x^2+y^2-2=0"

Hence, we are done.